A leopard is sleeping on a branch in a tree. While asleep, the leopard rolls off her branch
and falls from rest, hitting the ground 1.1 s later. How high is the branch from the ground?
1
Expert's answer
2020-12-14T12:17:57-0500

We can find the height of the branch from the kinematic equation:


"y=\dfrac{1}{2}gt^2=\dfrac{1}{2}\cdot 9.8\ \dfrac{m}{s^2}\cdot (1.1\ s)^2=6.0\ m."

Answer:

"y=6.0\ m."